A metal M form two different chloride. If 12.7g of chloride a contain 7.1g of chloride and 16.3g of chloride B contain 10.7g of chloride, show that the figures agree with the law of multiple proportional. Write their formula
1
Expert's answer
2020-02-03T07:46:16-0500

chloride A: MClx

chloride B: MCly


m - molar mass of M


w(Cl in A)=7.1/12.7=0.5591

w(Cl in B)=10.7/16.3=0.6564


0.5591=35.5x/(35.5x+m)

0.6564=35.5y/(35.5y+m)


19.85x+0.5591m=35.5x

23.30y+0.6564m=35.5y


0.5591m=15.65x

0.6564m=12.2y


m=28x

m=18.6y


x=2

y=3

m=56 (Fe)


Answer: A - FeCl2, B - FeCl3